# ShaRP: A Novel Feature Importance Framework for Ranking

### v3

@@ 1.1. Summary of Our Approach
As our final example that illustrates two other QoIs supported by ShaRP , consider Figure 3a that presents a summary view of feature contributions to the top-$k$ in CS Rankings. In this figure, the importance of each of the four features is summarized for the first 10% of the departments, followed by the next 10%, and so on. Perhaps unexpectedly, we observe that a strong publication record in systems is the most important feature for placing a CS department among the top-10, significantly ahead of AI.
Figures 3b3d analyzes the relative order among three departments that appear among the top-10: Georgia Tech (rank 5), Stanford (rank 6) and UMich (rank 7). We use visualizations for the pairwise QoI to explain that Georgia Tech ranks better than Stanford because of its relative strength in systems, and, further, that Stanford ranks better than UMich despite Stanford’s relative weakness in systems.
@@
(c) Pairwise QoI: Georgia Tech ranks higher than Stanford because of its relative strength in Systems.
@@
(d) Pairwise QoI: Stanford ranks higher than UMich despite Stanford’s relative weakness in Systems.
@@
Figure 3. Feature importance for the top-$k$ QoI (i.e., selection) for CS Rankings in 3a, with further analysis of the relative orders among two pairs of departments in 3c and 3d.
1.2. Contributions
Our first contribution is a formalization of several natural quantities of interest (QoI) for ranking, expressing feature contributions to an item’s score or rank, to its presence or absence at the top-$k$, and to the relative order between a pair of items. To define the QoIs, we build on the conceptual Quantitative Input Influence (QII) framework by Datta et al. (2016) and extend it to quantify the influence of features on the ranked outcome. QII quantifies feature importance based on Shapley values ((Shapley et al., 1953)), a canonical method of dividing revenue in cooperative games. Like other Shapley value methods, QII explains the influence of an item’s features (players) on the classification outcome for that item (cooperative game). Unlike other Shapley value methods, QII does not necessarily use the score of the model being explained as the payoff but allows for the definition of a variety of payoff functions, termed Quantities of Interest (QoIs). This flexibility is essential for explaining ranking, because of the need to capture the lack of independence between per-item outcomes.
QII is a conceptual framework and only a preliminary implementation is publicly available. Our second contribution is a robust and extensible open-source ShaRP  library that implements QII, making this powerful framework — including also its functionality for explaining classification outcomes — available to the community. ShaRP  is available as an open-source library at https://github.com/DataResponsibly/ShaRP and we already use it to teach interpretability to computer science and data science students. Our implementation includes approximate QoI computation to achieve interactive performance while retaining fidelity and consistency.
Our third contribution is an extensive experimental evaluation of ShaRP . We define metrics for evaluating ranking explanability methods, and present results of qualitative analysis of ShaRP  with numerous real and synthetic datasets, showing that feature importance for all QoIs depends on the properties of the data distribution even when the scoring function remains fixed. Further, we show that the QoIs we define bring interesting — and complementary — insights about the data that cannot be captured simply by the relationship between the scores and the ranks. We also present results of a large-scale quantitative evaluation of ShaRP , demonstrating the scalability of approximations, and comparing it to other local feature importance methods.
2. Related Work
Local feature-based explanations
Ribeiro et al. (2016b) proposed LIME, a feature-based method that explains a complex classification model locally using a class of more interpretable models such as sparse linear models.  Lundberg and Lee (2017) proposed SHAP, a feature-based method that uses Shapley values to explain a large variety of classification and regression models, and that generalizes LIME. Both LIME and SHAP have been implemented in broadly used software libraries, and they can be used to explain the score of an item (score QoI in our terminology). We compare with both SHAP and LIME experimentally in Section 7.4.
Feature-based explanations for ranking
Yang et al. (2018) proposed a “nutritional label” for score-based rankers that includes two visual widgets that explain the global ranking methodology and outcome: “Recipe” shows scoring feature weights and “Ingredients” shows the features (scoring or not) that have the strongest Spearman’s rank correlation with the score. The authors observed that the weight of a feature in the “Recipe” rarely corresponds to how well it correlates with the score. This work generates global feature-based explanations, showing which features matter more or less for the ranked outcome for all items. Another very recent line of work, Heuss et al. (2024), adapts SHAP (Lundberg and Lee, 2017) for learning-to-rank, to compute global feature-based explanations. In contrast, we focus on local explanations that show feature importance for the outcome of a specific item or a pair of items.
Gale and Marian (2020) developed “participation metrics” for score-based rankers that measure the contributions of scoring features to whether an item is among the top-$k$. Their most important metric, “weighted participation,” attributes an item’s presence among the top-$k$ to the values of the scoring features and their scoring function weights, to the distance of a feature value from the maximum possible values, and to whether the value exceeds the $k^{th}$ highest value for the feature. This quantity is then aggregated across all items at the top-$k$, and outputs one participation value per feature. The main difference between our method and Gale and Marian (2020) is that we generate local — per item — explanations, while their explanations are group-wise, for a specific group — the top-$k$. Our top-$k$ QoI is similar to their weighted participation, however, our focus is on local — per item — explanations, which we can then (optionally) aggregate across the top-$k$. As a result, we cannot compare with Gale and Marian (2020) directly.
Yuan and Dasgupta (2023) developed an interface for sensitivity analysis of ranked synthetic datasets with two normally distributed features, and with a scoring function with linear feature weights. They quantify the importance of each feature by subtracting its value from the mean value of that feature. Given their distributional assumptions and assumptions about the scoring function, this corresponds to an approximation of Shapley values of the features (Lundberg and Lee, 2017). They also propose two alternative methods to quantify feature contributions, calculated using linear weights and the score/rank of the item. We re-implemented and substantially extend their method to work with more than 2 features with arbitrary feature distributions, and with arbitrary scoring functions. This allowed us to compare with their method experimentally in Section 7.4.
Anahideh and Mohabbati-Kalejahi (2022) provide local feature-based explanations using multiple methods, including Shapley values computed with SHAP (Lundberg and Lee, 2017), and apply these methods to the neighboring items ranked around the item being explained. They justify this approach by observing that feature importance changes when different size ranking neighborhoods are considered. Like Anahideh and Mohabbati-Kalejahi (2022), we find that feature importance differs significantly per rank stratum. However, countering the assumptions of  Anahideh and Mohabbati-Kalejahi (2022), we show that slight changes in feature values can move an item significantly in the ranks, far beyond the local neighborhood. We will conduct an experimental comparison with their method in Section 7.4.
Finally,  Hu et al. (2022) developed PrefSHAP, a method for explaining pairwise preference data from learned rankers. They calculate the Shapley value contributions for pairwise preferences on a new quantity they propose: They create artificial items from the pair of items simultaneously for each coalition, and then evaluate how the preference changes. While this quantity can be applied to any preferential function, they introduce PrefSHAP specifically for functions that use the generalized preferential kernel defined in their own prior work (Chau et al., 2022) for running time efficiency. They compare their results with an adaptation of SHAP for preference models and show that this adaptation does not capture the contributions properly. We also argue that ranking-specific quantities of interest are needed to explain rankings and introduce a broader set of QoIs (score, rank, top-$k$, and pairwise). We do not conduct an experimental comparison between our pairwise QoI and PrefSHAP Hu et al. (2022) because our focus is on preferences that are induced by score-based rankers, while they work only with a specific kind of learned models.
In summary, we share motivation with these lines of work but take a leap, describing the first comprehensive Shapley-value-based framework for selection and ranking.
3. Preliminaries and Notation
Ranking.
A ranking $r_{\mathcal{D}}$ is a permutation over the items in $\mathcal{D}_{\mathit{}}$. Letting $n=\left|\mathcal{D}_{\mathit{}}\right|$, we denote by $r_{\mathcal{D}}=\langle\textbf{v}_{1},\ldots,\textbf{v}_{n}\rangle$ a ranking that places item $\textbf{v}_{i}$ at rank $i$. We denote by $r_{\mathcal{D}}(i)$ the item at rank $i$, and by $r_{\mathcal{D}}^{-1}(\textbf{v})$ the rank of item v in $r_{\mathcal{D}}$. In score-based ranking, we are interested in rankings induced by some scoring function $f$. We denote these rankings $r_{\mathcal{D},f}$. For example, in Figure 1(b), $r_{\mathcal{D},f}(1)=\texttt{Bob}$, $r_{\mathcal{D},f}^{-1}(\texttt{Leo})=7$. We assume that $r_{\mathcal{D},f}^{-1}(\textbf{v}_{1})<r_{\mathcal{D},f}^{-1}(\textbf{v}_{2})<\dots<r_{\mathcal{D},f}^{-1}(\textbf{v}_{n})$, where smaller rank means better position in the ranking.
@@ 3. Preliminaries and Notation
Shapley values
We will use Shapley values to explain ranked outcomes using the set of features $\mathcal{A}$ as the players, and the outcome (or the quantity of interest, QoI) as the payoff function. In addition to the definition of players and the payoff function, Shapley values require the quantification of the payoff over a subset of the players. This, in turn, requires some notion of estimating the payoff over a subset of the features. Consequently, for any Shapley value implementation, a method of feature removal or masking is required (Covert et al., 2021).
A common method (e.g., used in SHAP (Lundberg and Lee, 2017; Covert et al., 2021)), for a coalition (subset of features) $\mathcal{S}\subseteq\mathcal{A}$, is to marginalize out the features not in the coalition $\mathcal{A}\setminus\mathcal{S}$ and draw values from the marginal distributions of the subset of features in $\mathcal{S}$ jointly. Another alternative (e.g., used in QII (Datta et al., 2016)) is to draw values of each feature in $\mathcal{S}$ independently from its marginal distribution.
More specifically, let $\textbf{v}_{\mathcal{S}}$ denote a projection of v onto $\mathcal{S}$. In the example in Figure 1, $(\texttt{Bob},\texttt{4},\texttt{5},\texttt{2})_{\{name,gpa\}}=(\texttt{Bob},\texttt{4})$. We define a random variable U that draws values from the marginal distributions of the subset of features in $\mathcal{S}$. Let U $=\langle\textbf{u}_{1},\dots,\textbf{u}_{m}\rangle$ denote a vector of $m$ items sampled from $\mathcal{D}$ using this method. For a subset of features $\mathcal{S}\in\mathcal{A}$, let $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}=\langle\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}(\textbf{u}_{1})_{\mathcal{S}},\dots,\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}(\textbf{u}_{m})_{\mathcal{S}}\rangle$ denote a vector of items, in which each $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}(\textbf{u}_{i})_{\mathcal{S}}$ takes on the values of the features in $\mathcal{S}$ from $\textbf{u}_{i}$, and the values of the remaining features $\mathcal{A}\setminus\mathcal{S}$ from v. We calculate Shapley values using this set of features $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}$, note that if $m=|\mathcal{D}-1|$ we use the entire dataset $\mathcal{D}\setminus\textbf{v}$ to calculate the exact Shapley values.
Shapley values satisfy several natural axioms, including efficiency, symmetry, dummy, and additivity (Shapley et al., 1953), with additional useful properties, such as monotonicity, following from these axioms(Lundberg and Lee, 2017). Efficiency states that the sum of the contributions of all features for item v equals the QoI: $\sum_{i\in\mathcal{A}}\varphi_{i}(f,\textbf{v})=f(\textbf{v})$. Using this property, a Shapley-value-based explanation can be used to reconstruct the outcome, potentially leading to the explanation being more actionable, as we argued in Example 2. We will use the efficiency property to define the fidelity metric for comparing explanations (Section 6).
4. Quantities of Interest for Ranking
The first contribution of our work is that we define QoIs that are appropriate for ranked outcomes. In addition to the expected score, we also introduce rank, top-$k$, and pairwise QoI.
Score QoI
The Shapley value function for the score QoI is:
This QoI captures the impact of an item’s features on its score. This is the QoI used by the popular feature-based explanation methods such as SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016a). To get the contribution of a set of features $\mathcal{A}\setminus\mathcal{S}$, we take the expected value of the score over a random variable $\textbf{U}_{\mathcal{S}}$ that draws values from the marginal distributions of the set of features in $\mathcal{S}$.
Rank QoI
The Shapley value payoff function for the rank QoI is:
where $\mathcal{D}^{\prime}$ is $\mathcal{D}\cup\{\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}\}\setminus\textbf{v}$. This QoI evaluates the impact of an item’s features on its rank. To get the contribution of a set of features $\mathcal{A}\setminus\mathcal{S}$, we take the expected value of the rank over a random variable $\textbf{U}_{\mathcal{S}}$ that draws values from the marginal distributions of the set of features in $\mathcal{S}$.
Top-$k$ QoI
(4) $QoI_{f,\textbf{v},\mathcal{D}}(S)=\mathbb{E}_{\textbf{U}_{\mathcal{S}}}[\mathbbm{1}_{r_{\mathcal{D}^{\prime},f}(1\ldots k)}(\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}})]$
where $\mathcal{D}^{\prime}$ is $\mathcal{D}\cup\{\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}\}\setminus\textbf{v}$. The difference with rank QoI (Equation 3) is that here we compute the expectation over the indicator function that returns 1 if an items’ rank is at most $k$ and 0 otherwise.
Pairwise QoI
The Shapley value payoff function for the pairwise explanation of the difference in the ranks between items v and u is defined on $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{u}_{\mathcal{S}}$ instead of $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}$:
(5) $QoI_{f,\textbf{v},\mathcal{D}}(S)=r^{-1}_{\mathcal{D}^{\prime},f}(\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{u}_{\mathcal{S}})$
Using this function, we quantify the contribution of each feature of v to the relative order (preference) between the two items.
5. Implementation of ShaRP

### v4

@@ 4. Quantities of Interest for Ranking
Similarly, Figure 4a presents the aggregated feature contributions to the top-$k$ in CS Rankings. In this figure, the contribution of each of the four features to the top-$k$ is summarized for the highest-ranking 10% of the departments, followed by the next 10%, and so on. We again observe that a strong publication record in systems is the most important feature for placing a CS department among the top-10%, significantly ahead of AI. The feature importance it shows is consistent with Figure 3b (score QoI), but the high positive impact of Systems on a department’s presence at the top-$k$ is even stronger pronounced. Additionally, unlike the score QoI, this top-$k$ QoI shows that Theory is impactful for getting into the top-$k$.
Figure 2, previewed in the Introduction, presents a feature-based explanation for Texas A&M in CS Rankings, where it has a score of 3.941 and is ranked in position 34. Figures 2a and 2b show feature importance for the score and rank QoIs, respectively. Both are presented as waterfall plots (Lundberg and Lee, 2017). This plot displays how each feature contributed to the difference of the item’s outcome $f(x)$ from the mean outcome $\mathbb{E}[f(X)]$. For example, Texas A&M’s 0.16 AI value helped it score 0.45 higher than the mean score of 2.722. Features in red in that figure positively impact the QoI and in blue negatively. The sign depends on whether a higher (score) or a lower (rank) QoI is better. The sum of feature importance values equals the difference between the item’s QoI and the mean. Finally, the magnitude of feature importance depends on the QoI and is hence not directly comparable across QoIs.
According to Figure 2a, “Interdisciplinary” is the most (positively) impactful feature on the score of Texas A&M, followed by “AI” and then by “Systems”, while “Theory” contributes negatively. In contrast, all features contribute positively to the rank of Texas A&M, with “Systems” being the most impactful. This observation illustrates that different QoI are useful for answering different questions. If the goal is to improve the score, then Texas A&M should invest in interdisciplinary and AI research. However, if the goal is to improve the rank, then it’s better to prioritize systems research. Why are the score and rank QoI explanations different? The reason is that a change in score may not result in a change in rank. Specifically, for a rank to change, the change in the score must be large enough to surpass the score of the item ranked directly above.
Another key aspect of these plots is the color of each feature, which indicates whether a feature contributes positively or negatively to the outcome. The contributions add up to the difference from the mean outcome so, their magnitude and color vary depending on the QoI. For example, the average score is influenced by outliers, while rank is not (by construction), so we expect the color of the contributions to be very different. Indeed, in CS Rankings, over 70% of departments have scores below the mean. As a result, when using the score QoI, many or all of their features appear to contribute negatively. This highlights that the meaning of positive and negative contributions is dependent on the chosen QoI. See Figure 9b in the Appendix for the score vs. rank distribution for this dataset.
@@
(c) Pairwise method using the rank QoI: Georgia Tech ranks higher than Stanford because of its relative strength in Systems.
@@
(d) Pairwise method using the rank QoI: Stanford ranks higher than UMich despite Stanford’s relative weakness in Systems.
@@
Figure 4. Feature importance for the top-$k$ QoI (i.e., selection) for CS Rankings in 4a, with further analysis of the relative orders among two pairs of departments in 4c and 4d.
5. Shapley value Method for Pairwise Explanations
@@
In Section 3 we discussed three different feature removal techniques for Shapley values. Here, we introduce a new Shapley value explanation method that uses the baseline feature removal technique. Note that, as we explained in Section 4, all QoIs can be used with any feature removal method so, any QoI can be used to calculate the pairwise explanation.
Baseline feature removal techniques select one “neutral” item as the “baseline” and then compare all other items to it. The benefit of these methods is that the exact feature contributions can be computed without any sampling. The disadvantage is that it is often hard to select the baseline sample because different baseline samples create different feature attributions and, in most contexts, it is hard to identify a “neutral” or “average” item. As an example, in related work, we mentioned  (Singh and Anand, 6 29) that attempts to identify a good baseline input document for DeepSHAP in IR. As another example, the baseline implementation of SHAP (Lundberg and Lee, 2017) uses the all-zeroes item as the baseline sample. While selecting a baseline sample is not simple in most cases, we find that the baseline feature removal technique is a natural fit when we are explaining the difference in outcomes between two items v and u.
When explaining the pairwise outcome of two items v and u, we are going to generate an explanation for one item using the other as the baseline. In other words, for coalition $\mathcal{S}$, we will be creating the hybrid item $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{u}_{\mathcal{S}}$. Note that we do not need the feature distributions or any other parameters for this method. Additionally, note that we are not selecting a fixed item as the baseline but we dynamically change it depending on the pair of items we want to compare. This definition has a natural interpretation, the feature importance of a pairwise explanation amounts to the difference between the outcome of the two items. According to the property of efficiency (see Section 3) we have: $\sum_{i\in\mathcal{A}}\varphi_{i}(f,\textbf{v})=f(\textbf{v})-\mathbb{E}_{\textbf{X}}[f(\textbf{X})]=f(\textbf{v})-\frac{1}{2}(f(\textbf{v})+f(\textbf{u}))=\frac{1}{2}(f(\textbf{v})-f(\textbf{u}))$.
The Shapley value of v in comparison to u is defined as:
(7) $\phi_{i}(f,\mathbf{v}\succ\textbf{u})=\sum_{\mathcal{S}}\frac{|\mathcal{S}|!(n-|\mathcal{S}|-1)!}{n!}\iota_{f,\mathbf{v},\textbf{u}}(i,\mathcal{S})$
Note that Eq. 7 differs from Eq. 6 in setting $\mathcal{D}=\{\textbf{u}\}$. Hence, in Section 7 we are able to implement both the baseline and the marginal feature removal techniques using one algorithm. Because pairwise preferences are of special interest to ranking tasks, we will only be using rank as the QoI for the pairwise method in the rest of the paper.
Case Study: Explanations of Pairwise Outcomes for CSRankings
In Figure 4b4d we continue our analysis of the top-$k$ and consider the relative ranking of three universities: Georgia Tech in rank 5, Stanford in rank 6, and UMich in rank 7. We use pairwise Shapley explanations that can clarify rank differences between two items. Consider Fig. 4c, where we explain the pairwise outcome for Georgia Tech $\textbf{v}_{\text{Georgia Tech}}=(28.5,7.8,6.9,10.2)$ vs $\textbf{u}_{\text{Stanford}}=(36.7,5.4,13.3,11.5)$. To explain the pairwise outcome difference, for each coalition, we construct $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{u}_{\mathcal{S}}$. For example, for coalition $\mathcal{S}=\{\text{AI},\text{Systems}\}$, we construct $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{u}_{\mathcal{S}}=(36.7,5.4,6.9,10.2)$, enabling a direct feature comparison of the item being explained ($\textbf{v}_{\text{Georgia Tech}}$) against the baseline sample ($\textbf{u}_{\text{Stanford}}$). We wish to understand why Georgia Tech is ranked higher than Stanford (Figure 4c), and why Stanford is ranked higher than UMich (Figure 4d). In both cases, Georgia Tech and UMich have lower values than Stanford for all features except Systems. In Figure 4c, we see that indeed all features except Systems contribute negatively in the relative rank of Georgia Tech as opposed to Stanford. However, the Systems value of Georgia Tech is high enough to rank it higher than Stanford. Reversely, for UMich, we see that, while Systems is the most important feature in the top-10% stratum (Figure 4a), Stanford is ranked higher because of the aggregate contributions of the other features.
The pairwise explanation from ShaRP  is intuitive: for example, Systems improves Georgia Tech’s rank by 5 compared to Stanford. in other words, feature contributions sum to (half) the rank difference between these universities, aligning with Fig. 4a, which highlights Systems as particularly influential for top-$k$ universities.
6. Evaluating ranking explanation methods
We define several metrics for evaluating feature importance methods for ranking. Our focus is on explanation methods that return a numerical vector of feature attributions explaining the outcome for a given item. We can assess pair-wise explanation agreement by comparing the feature vectors of a pair of explanations. Furthermore, we can use an explanation to compute the outcome for the item being explained (e.g., its rank), and compare it to the actual observed outcome for that item. This allows us to assess fidelity of an explanation. Below, we describe explanation agreement and fidelity metrics and also explain how these primitives can be aggregated to assess sensitivity and fidelity of an explanation method, and to quantify inter-method explanation agreement.
Notation
In Section 3, we have been using $\phi(\textbf{v})$ to represent the vector of feature weights, computed using Shapley values. We generalize our definition here to $g(\textbf{v})$ to represent the output of any feature-based explanation method $g$, regardless of whether it consists of Shapley values or of some other numerical quantification of feature importance. For all methods we consider, $g(\textbf{v})$ is a vector of numerical contributions of each feature towards the outcome for item v.
6.1. Fidelity Metrics
Explanation Fidelity
A useful property of feature-based explanations is that the actual outcome can be computed from them. For Shapley-value-based explanations, this follows from the efficiency property of Shapley values, see Section 3. Fidelity measures how well the explanation $g(\textbf{v})$ matches the model prediction $f(\textbf{v})$ being explained, see (Chowdhury et al., 2023; Molnar, 2020). SHAP and LIME explanations can be used to compute an item’s score (score QoI in our terminology), with feature importance indicating the displacement due to that feature from the mean score, either positively or negatively. ShaRP  explanations can be used to compute the outcome for all supported QoIs, including score, rank, and top-$k$, and for the pairwise method.
For the QoIs, namely, score, rank, and top-$k$, we compute fidelity of explanation $g$ for item v as:
Here, $\mathsf{QoI}(\textbf{v})$ returns the value of the quantity of interest (i.e., the outcome being explained by $g$), such as v’s score, rank, or presence in the top-$k$, while $g(i,\textbf{v})$ is the contribution of the i-th feature of $v$. Finally, $Z$ is the normalizer set to the maximum distance between a pair of outcomes for the given dataset $\mathcal{D}$ and ranker $f$ (omitted to simplify notation), and for the specified QoI. Note that, for pairwise explanations, fidelity $F(g,\textbf{u}\succ\textbf{v})=1$ if $u$ is ranked higher than $u$ and if $g$ predicts that relative order among the items, and is 00 otherwise.
Example 0.
Method Fidelity
For pairwise, we compute $F(g,\mathcal{D})$ as the expectation of $F(g,\textbf{u}\succ\textbf{v})$ over all pairs of distinct items $\textbf{u},\textbf{v}\in\mathcal{D}$.
6.2. Agreement Metrics
Explanation Agreement
[…]
Example 0.
[…]
Method Agreement
Method Sensitivity
7. The  ShaRP  library